Baichung's father is 26 years younger than Baichung's grandfather and 29 years older than his son.If the sum of their ages is 135 than what is the age of all the 3? Please send explanation too.
Answer by ankor@dixie-net.com(22740)

Baichung's father is 26 years younger than Baichung's grandfather and 29 years older than his son.
If the sum of their ages is 135 than what is the age of all the 3?
:
let b = B's age
th
b + 26 = grandfathers age
and
b - 29 = son's age
:
Their sum
b + (b+26) + (b-29) = 135
:
Solve for b, then you can find that ages of all three